# When Large Language Models Meet Vector Databases: A Survey

### v3

@@ Abstract
Index Terms:
Large Language Models, Retrieval-Augmented Generation, Vector Databases
I Introduction
The field of Natural Language Processing (NLP) achieves a great milestone when ChatGPT unlocks the power of interactive smart assistants and people are offered the unprecedentedly tangible experience of talking to a machine program that is seemingly more knowledgeable than an ordinary person. This remarkable breakthrough is largely attributed to the advancements in LLMs, such as GPT  [1, 2], T5  [3], and Llama  [4]. They can process, understand, and generate human-like text. Thanks to the power of pre-training, which involves training a language model on a massive corpus of text data, LLMs can capture the complexities of human languages, including context, idioms, and even cultural references.
However, despite the impressive capabilities of LLMs, they are still under the shadows of doubt in certain aspects. 1. One major shortcoming is the problem of hallucination, where LLMs generate plausible but factually incorrect or faithfully nonsensical information  [5]. There are many potential causes behind this. First, LLMs lack domain knowledge. The fact that LLMs are primarily trained on public datasets  [6] inevitably leads to a limited ability to answer domain-specific questions that are out of the scope of their internal knowledge. Moreover, real-time knowledge updates are challenging for LLMs. Even if the questions are within the learning corpus of LLMs, their answers may still exhibit limitations because the internal knowledge may be outdated when the outside world is dynamic and keeps changing  [7]. Last but not least, LLMs are found to be biased. The datasets used to train LLMs are large, which may introduce systematic errors  [8]. Essentially, every dataset can be questioned with bias issues, including imitative falsehoods  [9], duplicating biases  [10], and social biases  [11]. 2. Another disadvantage of incorporating LLMs commercially is the expensive cost [12]. For an average business entity, applying LLMs for business use is barely feasible. It is almost impossible for a non-tech company to customize and train a GPT model of its own because they don’t have the resources and talents to conduct such a big project  [13], while frequent API calls to third-party LLM providers like OpenAI can be extremely expensive, not to mention there are a very limited number of such providers in certain areas. 3. The oblivion problem of LLMs has been controversial because LLMs are found to tend to forget information from previous inputs. It is studied that LLMs also exhibit the behavior of catastrophic forgetting  [14] as neural networks do  [15].
@@ I Introduction
Since there are lacking papers that introduce LLMs in the view of VecDB, this survey aims to picture how VecDBs can be potential solutions to refine LLMs’ known shortcomings in previous works and hopes to offer a unique perspective of future directions in the intersection that is fertile of research opportunities.
TABLE I: Comparison of mainstream all-level databases supporting vector data (the version number represents the first version that supports vector search and its release date).
Abbreviations note: Ftx., Rel., Vec. stand for full-text search, relational database, and vector search.
@@
II Background
II-A Large Language Models
@@
II-A1 Development of Language Models
@@
With the advent of neural networks, the field of NLP has undergone a transformative shift, which began with the introduction of Recurrent Neural Networks (RNNs)  [16]. RNNs provide a way to process sequences of words and capture temporal dependencies in text. And 2 of its famous variants, Gated Recurrent Units (GRUs)  [17] and Long Short-Term Memory networks (LSTMs)  [18] address the limitations of RNNs in handling problems of vanishing and exploding gradients.
@@
The next pivotal milestone for language models is the widespread adoption of the Transformer architecture  [19] which has set a new standard for language models. The multi-head self-attention modules and cross-attention modules in the encoders and decoders enable the model to capture long-range dependencies, parallel processing, and contextual understanding. Furthermore, the model starts to rapidly grow in size in the Transformer era. More importantly, the Transformer represents a paradigm shift in the NLP field, and the success of the Transformer architecture becomes the significant backbone of LLMs.
II-A2 Are We There Yet? Challenges Faced by Pure LLMs
@@
Although LLMs have achieved remarkable success, they face significant challenges that cannot be overlooked. Addressing these challenges is crucial for the continued advancement and practical application of these models.
@@
One major challenge is hallucinations in LLMs  [20], which undermine the reliability of their outputs. LLMs often generate plausible yet factually incorrect or nonsensical information  [5]. Several factors contribute to this issue. First, LLMs lack domain-specific knowledge, primarily due to their training on public datasets  [6], limiting their ability to answer domain-specific questions. Additionally, updating LLMs with real-time knowledge is a challenge, leading to outdated internal knowledge in a rapidly changing world  [7].
Another notable concern is the substantial computational resources required for training and operating LLMs, which raises environmental and accessibility concerns  [21] and underscores the importance of sustainable AI practices.
@@
Transitioning to VecDBs could potentially address these challenges. VecDBs, with their efficient data structuring and retrieval capabilities, offer a promising solution for enhancing model editing flexibility, reducing hallucinations by improving domain-specific responses, and optimizing computational resource utilization, thereby aligning more closely with the principles of sustainable AI.
II-B Vector Databases, the V-factor
II-B1 A distinct database optimized for vector data
@@
While LLMs like ChatGPT are relatively new, database management systems (DBMS) have been thoroughly developed and applied in many aspects for the last 60 years. DBMS are well-recognized for their consistent stability and universality for structured data with fixed formats that excel well with computer hierarchical computing architecture and storage. However, the development and wide application of deep learning models such as convolutional neural networks [22] and transformers [23], enable the embedding of unstructured multi-modal data like images and text mapped into corresponding fixed-length vector representations, which include the high-dimensional semantic features of original data, and the semantic similarities are naturally represented by distances between vectors, requiring a new type of DBMS that is specifically designed for handling vector data operations, especially search and storage.
@@
II-B2 Efficient vector retrieval within VDBMS
@@
Since brute-force search is time-consuming and computationally expensive, this calls for a more efficient search and storage technique with satisfying accuracy. Vector indexing can solve both of these problems, which optimized for ANN search within VDBMS include tree-based [24, 25], hash-based [26], and graph-based methods [27]. At the same time, the ANN Benchmarks  [^1] https://ann-benchmarks.com/ has also showcased the great performance gap between brute-force search and index-enhanced ANN search, as well as those indexing techniques combining different vector quantization methods such as Product Quantization (PQ) [28]. All these operations could be efficiently done based on the vector indexes, which are optimally designed collections of vectors deployed together for approximate nearest neighbor (ANN) search. The ANN Benchmarks have also showcased the great performance gap between brute-force search and index-enhanced ANN search. Also, according to them, among all these indexes, the graph-based Hierarchical Navigable Small Worlds (HNSW) provides state-of-the-art performance and capability with great universality and is widely used within most VecDB management systems, including the examples we mentioned in Table I.
II-B3 A feasible solution of knowledge base for LLMs
@@
Traditional full-text and keyword search engines, such as Elastic Search and Amazon OpenSearch, are based on term frequency metrics like BM25, which has proved practical for many search applications. However, such search techniques require significant investment in time and expertise to tune them to account for the meaning or relevance of the terms searched, you can’t search it like you mean it. On the other hand, they lack multi-modal storage and retrieval support, as VecDBs do.
@@
Nevertheless, VecDBs are capable of solving the aforementioned problem by simply transforming data into vectors and leveraging their retrieval efficiency; thus, the emergence of VecDBs (as shown in Table I) has greatly influenced machine learning systems and their multi-modal applications. An intuitive application using vector search is searching images with an image on search engines built on VecDBs [29] like Google Image Search  [^2] https://cloud.google.com/vertex-ai/docs/vector-search/overview, which uses one input image to find similar images on the internet. While VecDBs boast their unique search capability and efficiency for unstructured data, they are just a back-end factor for manifold applications. To maximize their abilities, this leads us to an interesting question: since LLMs use embeddings to represent text as vectors, can developers combine VecDBs and LLMs to overcome the aforementioned challenges that are inherited in LLM applications? The answer is yes.
@@
As a kind of database encapsulated with vector search in vector data that represents real-world information within high dimensionalities, VecDBs are well-capable for retrieval applications [30] incorporating LLMs because LLM applications are generally read-intensive, not requiring many write-related changes, especially data deletes [31]. On the other hand, VecDBs can efficiently manage and warehouse vector data required and generated by LLMs, thus providing a solid data cornerstone for both LLMs and their applications. While LLMs are often limited by domain knowledge that cannot be uploaded or distributed due to security and privacy concerns, domain knowledge is often unstructured data that can easily be embedded into VecDBs for efficient local retrieval and further integration with generative AIs. Moreover, the computing and storage resources of VecDBs are way cheaper than LLMs, since they do require costly GPUs and TPUs, thus achieving a cost-effective way of fast retrieval and durable (non-volatile) storage.
III Synergizing LLMs with VecDBs: Enhancements and Innovations
@@
III-A VecDB as an External Knowledge Base: Retrieval-Augmented Generation (RAG)
III-A1 Development of RAG Paradigm
@@
Data storage means establishing reliable external memory like VecDBs, and there are detailed recipes for this process [32]. It starts with data preprocessing, during which the original data is collected, cleaned, integrated, transformed, normalized, and standardized. The processed data is chunked into smaller segments because of the contextual limitations of LLMs. These segments of data are then converted by an embedding model into vectors, the semantic representations of the data, which are stored in VecDBs and will be used for vector search in later steps. A well-developed VecDB will properly index the data and optimize the retrieval process. The retrieval part starts with a user asking a question in the form of prompts to the same embedding model, which has generated the vector representation of the stored data and gained the vector embeddings of the question. The next step of the process is vector searching and scoring inside the VecDBs, which essentially involves computing similarity scores among vectors, and the database then identifies and retrieves the data segments that have the highest similarity scores (top $K$ in most RAG systems) compared to the query vector. These retrieved segments are then converted back from their vector format to their original form, which is typically the text of documents. In the generation part, LLMs are involved in generating the final answers. The retrieved documents, along with the user’s question, are incorporated into a specifically chosen and designed prompt template. This template selection is based on the task type and aims to effectively merge the context with the question to form a coherent prompt. The selected LLM is provided with the prompt and generates the final answer.
III-A2 RAG with VecDBs
In the prototype of the RAG paradigm, we use such an idea to overcome the hallucination problems of LLMs by providing domain-specific information with accurate instructions like Chain-of-Thought [33], where VecDBs serve as an external knowledge base to warehouse domain-specific related information, then LLMs can easily reason with question-related information from massive data owned by users.
@@ III-A2 RAG with VecDBs
Many research and industry examples justified RAG’s profound effect. Azure AI Search (formerly Microsoft Cognitive Search) developed its vector search using Qdrant  [^3] https://learn.microsoft.com/en-us/azure/search/vector-search-overview. Another example is Pinecone’s Canopy  [^4] https://www.pinecone.io/blog/canopy-rag-framework/, an open-source framework that leverages Pinecone’s VecDB to build and host production-ready RAG systems. Companies like Spotify and Yahoo have adopted Vespa, an AI-driven online VecDB, and Yahoo uses it to help users directly chat with their mailbox, ask questions about their mailbox, and tell it to do things for them. This uses personalized data integrated with LLMs to form a RAG system  [^5] https://vespa.ai/. These examples demonstrate the strength of RAG combined with VecDBs to address the unique challenges faced by business enterprises in managing and extracting value from diverse datasets but also showcase how the aforementioned difficulties faced by LLMs can be solved via integration with VecDBs.
III-B VecDB as a Cost-effective Semantic Cache
The combination of VecDBs and LLM not only facilitates the profound application of LLM with RAG but also provides a new frontier for cost-effective end-to-end LLM applications [34].

### v4

@@ Abstract
1 Introduction
The field of Natural Language Processing (NLP) achieves a great milestone when ChatGPT unlocks the power of interactive smart assistants and people are offered the unprecedentedly tangible experience of talking to a machine program that is seemingly more knowledgeable than an ordinary person. This remarkable breakthrough is largely attributed to the advancements in LLMs, such as GPT  Brown et al. (2020); Achiam et al. (2023), T5  Raffel et al. (2020), and Llama  Touvron et al. (2023). They can process, understand, and generate human-like text. Thanks to the power of pre-training, which involves training a language model on a massive corpus of text data, LLMs can capture the complexities of human languages, including context, idioms, and even cultural references.
However, despite the impressive capabilities of LLMs, they are still under the shadows of doubt in certain aspects. 1. One major shortcoming is the problem of hallucination, where LLMs generate plausible but factually incorrect or faithfully nonsensical information  Huang et al. (2023). There are many potential causes behind this. First, LLMs lack domain knowledge. The fact that LLMs are primarily trained on public datasets  Penedo et al. (2023) inevitably leads to a limited ability to answer domain-specific questions that are out of the scope of their internal knowledge. Moreover, real-time knowledge updates are challenging for LLMs. Even if the questions are within the learning corpus of LLMs, their answers may still exhibit limitations because the internal knowledge may be outdated when the outside world is dynamic and keeps changing  Onoe et al. (2022). Last but not least, LLMs are found to be biased. The datasets used to train LLMs are large, which may introduce systematic errors  Bender et al. (2021). Essentially, every dataset can be questioned with bias issues, including imitative falsehoods  Lin et al. (2022), duplicating biases  Lee et al. (2022), and social biases  Narayanan Venkit et al. (2023). 2. Another disadvantage of incorporating LLMs commercially is the expensive cost Schwartz et al. (2020). For an average business entity, applying LLMs for business use is barely feasible. It is almost impossible for a non-tech company to customize and train a GPT model of its own because they don’t have the resources and talents to conduct such a big project  Musser (2023), while frequent API calls to third-party LLM providers like OpenAI can be extremely expensive, not to mention there are a very limited number of such providers in certain areas. 3. The oblivion problem of LLMs has been controversial because LLMs are found to tend to forget information from previous inputs. It is studied that LLMs also exhibit the behavior of catastrophic forgetting  Luo et al. (2023) as neural networks do  Kemker et al. (2017).
@@ 1 Introduction
In the light of lacking papers that introduce LLMs in the view of VecDB, this survey aims to picture how VecDBs can be potential solutions to refine LLMs’ known shortcomings in previous works and hopes to offer a unique perspective of future directions in the intersection that is fertile of research opportunities.
@@
Table 1: Comparison of mainstream all-level databases supporting vector data (the version number represents the first version that supports vector search and its release date). Abbreviations note: Ftx., Rel., Vec. stand for full-text search, relational database, and vector search.
2 Background
2.1 Large Language Models
@@
2.1.1 Development of Language Models
@@
With the advent of neural networks, the field of NLP has undergone a transformative shift, which began with the introduction of Recurrent Neural Networks (RNNs)  Zaremba et al. (2014). RNNs provide a way to process sequences of words and capture temporal dependencies in text. And 2 of its famous variants, Gated Recurrent Units (GRUs)  Chung et al. (2014) and Long Short-Term Memory networks (LSTMs)  Shi et al. (2015) address the limitations of RNNs in handling problems of vanishing and exploding gradients.
@@
The next pivotal milestone for language models is the widespread adoption of the Transformer architecture  Vaswani et al. (2017) which has set a new standard for language models. The multi-head self-attention modules and cross-attention modules in the encoders and decoders enable the model to capture long-range dependencies, parallel processing, and contextual understanding. Furthermore, the model starts to rapidly grow in size in the Transformer era. More importantly, the Transformer represents a paradigm shift in the NLP field, and the success of the Transformer architecture becomes the significant backbone of LLMs.
2.1.2 Power of Pre-training
The scaling laws proposed by OpenAI  Kaplan et al. (2020) highlight a critical trend: the scaling of Pretrained Language Models (PLMs), in terms of both model size and data volume, leads to significant improvement in downstream tasks. This is evidenced by the development of increasingly larger PLMs, such as the 1.76-trillion-parameter GPT-4  Achiam et al. (2023) and the 540-billion-parameter PaLM  Chowdhery et al. (2023). Unlike their smaller predecessors, like the 330-million-parameter BERT  Devlin et al. (2018) and 1.5-billion-parameter GPT-2  Radford et al. (2019), these large-scale models demonstrate unique behaviors and emergent abilities  Wei et al. (2022) in various tasks like zero-shot learning, which is extremely challenging for the small-scale models.
The power of pre-training lies in its ability to provide models with a general understanding of the languages, which is then tailored through additional training, known as fine-tuning, for specific tasks. The advantage is that the model doesn’t start from scratch when learning a new task; it builds upon a rich, pre-existing foundation of language understanding. This enables what is known as transfer learning  Yosinski et al. (2014).
2.1.3 Are We There Yet? Challenges Faced by Pure LLMs
@@
Although LLMs have achieved remarkable success, they face significant challenges that cannot be overlooked. Three primary challenges are particularly noteworthy in the context of pure LLMs. Addressing these challenges is crucial for the continued advancement and practical application of these models. We have outlined these challenges as follows:
  • Model Editing involves integrating more flexible architectures into models, enabling targeted updates or edits without the need for complete retraining. This presents a complex challenge: alterations in one part of the model can unexpectedly affect other areas. Ensuring that these updates neither degrade performance nor introduce biases is a difficult task. Model editing also includes elements such as personality, emotions, opinions, and beliefs. While these aspects have received some attention, they largely remain unexplored territories  Yao et al. (2023). Ethical concerns, including data biases  Raffel et al. (2020), privacy issues  Li et al. (2023a), and the potential misuse of generative content  Ganguli et al. (2022), are significant challenges in model editing.
@@
  • Hallucinations in LLMs  Ji et al. (2023) pose a major challenge, undermining the reliability of their outputs. LLMs often generate plausible yet factually incorrect or nonsensical information  Huang et al. (2023). Several factors contribute to this issue. First, LLMs lack domain-specific knowledge, primarily due to their training on public datasets  Penedo et al. (2023), limiting their ability to answer domain-specific questions. Additionally, updating LLMs with real-time knowledge is a challenge, leading to outdated internal knowledge in a rapidly changing world  Onoe et al. (2022).
  • Substantial computational resources required for training and operating LLMs raise environmental and accessibility concerns  Strubell et al. (2019), contradicting the principles of sustainable AI  Wu and Raghavendra (2022). The escalating sizes of LLMs and their training datasets significantly influence training and inference costs. The environmental footprint of these technologies is substantial, with significant greenhouse gas emissions and energy consumption during model training and tuning  Kaack et al. (2022). This situation necessitates a critical evaluation of AI’s environmental impact, encouraging the development of more energy-efficient models and the adoption of green computing practices  Wynsberghe (2021).
@@
Transitioning to VecDBs could potentially comprehensively address these challenges. VecDBs, with their efficient data structuring and retrieval capabilities, offer a promising solution for enhancing model editing flexibility, reducing hallucinations by improving domain-specific responses, and optimizing computational resource utilization, thereby aligning more closely with the principles of sustainable AI.
2.2 Vector Databases, the V-factor
2.2.1 A distinct database optimized for vector data
@@
While LLMs like ChatGPT are relatively new concepts, database management systems (DBMS) have been thoroughly developed and applied in many aspects in the last 60 years of history and are well-recognized for their consistent stability and universality for structured data with fixed formats that excel well with computer storage. However, the development and wide application of deep learning models such as convolutional neural networks He et al. (2016) and transformers Devlin et al. (2018), enable the embedding of unstructured multi-modal data like images and text mapped into corresponding fixed-length vector representations, which include the high-dimensional semantic features of original data, and the semantic similarities are naturally represented by distances between vectors, requiring a new type of DBMS that is specifically designed for handling vector data operations, especially search and storage.
@@
2.2.2 Efficient vector retrieval within VDBMS
@@
Since brute-force search is time-consuming and computationally expensive, this calls for a more efficient search and storage technique with satisfying accuracy. Vector indexing can solve both of these problems, which optimized for ANN search within VDBMS include tree-based Muja and Lowe (2009); Tao et al. (2009), hash-based Andoni and Indyk (2008), Product Quantization (PQ) Jégou et al. (2011), and graph-based methods Malkov and Yashunin (2018), while the ANN Benchmarks  [^1] https://ann-benchmarks.com/ has also showcased the great performance gap between brute-force search and index-enhanced ANN search. All these operations could be efficiently done based on the vector indexes, which are optimally designed collections of vectors deployed together for approximate nearest neighbor (ANN) search. The ANN Benchmarks have also showcased the great performance gap between brute-force search and index-enhanced ANN search. Also, according to them, among all these indexes, the graph-based Hierarchical Navigable Small Worlds (HNSW) provides state-of-the-art performance and capability with great universality and is widely used within most VecDB management systems, including the examples we mentioned in Table 1.
2.2.3 A feasible solution of knowledge base for LLMs
@@
Traditional full-text and keyword search engines, such as Elastic Search and Amazon OpenSearch, are based on term frequency metrics like BM25, which has proved practical for many search applications. However, such search techniques require significant investment in time and expertise to tune them to account for the meaning or relevance of the terms searched. On the other hand, they lack multi-modal storage and retrieval support, as VecDBs do.
@@
Nevertheless, VecDBs are capable of solving the aforementioned problem by simply transforming data into vectors and leveraging their retrieval efficiency; thus, the emergence of VecDBs (as shown in Table 1) has greatly influenced machine learning systems and their multi-modal applications. An intuitive application using vector search is searching images with an image on search engines built on VecDBs Wang et al. (2021) like Google Image Search  [^2] https://cloud.google.com/vertex-ai/docs/vector-search/overview, which uses one input image to find similar images on the internet. While VecDBs boast their unique search capability and efficiency for unstructured data, they are just a back-end factor for manifold applications. To maximize their abilities, this leads us to an interesting question: since LLMs use embeddings to represent text as vectors, can developers combine VecDBs and LLMs to overcome the aforementioned challenges that are inherited in LLM applications? The answer is yes.
@@
As a kind of database encapsulated with vector search in vector data that represents real-world information within high dimensionalities, VecDBs are well-capable for retrieval applications Asai et al. (2023) incorporating LLMs because LLM applications are generally read-intensive, not requiring many write-related changes, especially data deletes Pan et al. (2023). On the other hand, VecDBs can efficiently manage and warehouse vector data required and generated by LLMs, thus providing a solid data cornerstone for both LLMs and their applications. While LLMs are often limited by domain knowledge that cannot be uploaded or distributed due to security and privacy concerns, domain knowledge is often unstructured data that can easily be embedded into VecDBs for efficient local retrieval and further integration with generative AIs. Moreover, the computing and storage resources of VecDBs are way cheaper than LLMs, since they do require costly GPUs and TPUs, thus achieving a cost-effective way of fast retrieval and durable (non-volatile) storage.
3 Synergizing LLMs with VecDBs: Enhancements and Innovations
@@
3.1 VecDB as an External Knowledge Base: Retrieval-Augmented Generation (RAG)
3.1.1 Development of RAG Paradigm
@@
Data storage means establishing reliable external memory like VecDBs, and there are detailed recipes for this process Han et al. (2023). It starts with data preprocessing, during which the original data is collected, cleaned, integrated, transformed, normalized, and standardized. The processed data is chunked into smaller segments because of the contextual limitations of LLMs. These segments of data are then converted by an embedding model into vectors, the semantic representations of the data, which are stored in VecDBs and will be used for vector search in later steps. A well-developed VecDB will properly index the data and optimize the retrieval process. The retrieval part starts with a user asking a question in the form of prompts to the same embedding model, which has generated the vector representation of the stored data and gained the vector embeddings of the question. The next step of the process is vector searching and scoring inside the VecDBs, which essentially involves computing similarity scores among vectors, and the database then identifies and retrieves the data segments that have the highest similarity scores (top $K$ in most RAG systems) compared to the query vector. These retrieved segments are then converted back from their vector format to their original form, which is typically the text of documents. In the generation part, LLMs are involved in generating the final answers. The retrieved documents, along with the user’s question, are incorporated into a specifically chosen and designed prompt template. This template selection is based on the task type and aims to effectively merge the context with the question to form a coherent prompt. The selected LLM is provided with the prompt and generates the final answer.
3.1.2 RAG with VecDBs
In the prototype of the RAG paradigm, we use such an idea to overcome the hallucination problems of LLMs by providing domain-specific data with accurate instructions. where VecDBs serve as an external knowledge base to warehouse domain-specific data, then LLMs can easily handle massive data owned by users.
@@ 3.1.2 RAG with VecDBs
Many research and industry examples justified RAG’s profound effect. Azure AI Search (formerly Cognitive Search) developed its vector search using Qdrant  [^3] https://learn.microsoft.com/en-us/azure/search/vector-search-overview. Another example is Pinecone’s Canopy  [^4] https://www.pinecone.io/blog/canopy-rag-framework/, an open-source framework that leverages Pinecone’s VecDB to build and host production-ready RAG systems. Companies like Spotify and Yahoo have adopted Vespa, an AI-driven online VecDB, and Yahoo uses it to help users directly chat with their mailbox, ask questions about their mailbox, and tell it to do things for them. This uses personalized data integrated with LLMs to form a RAG system  [^5] https://vespa.ai/. These examples demonstrate the strength of RAG combined with VecDBs to address the unique challenges faced by business enterprises in managing and extracting value from diverse datasets but also showcase how the aforementioned difficulties faced by LLMs can be solved via integration with VecDBs.
3.2 VecDB as a Cost-effective Semantic Cache
The combination of VecDBs and LLM not only facilitates the profound application of LLM with RAG but also provides a new frontier for cost-effective end-to-end LLM applications.